# The Learning Approach to Games

### v2

@@ 3 Stated Mean-Field Games
$p(t,x,\mu,a;\cdot)\text{ is a probability measure on $\mathbb{S}_{t}$, for all $t\in\mathbb{T},x\in\mathbb{S}_{t},\mu\in{\cal P}(\mathbb{S}_{t})$}$
@@
Note that $\mu^{\Xi}$ corresponds to (2.3) for the population. If the second marginal of $\Xi$ is a Dirac measure $\delta_{\alpha}$ independent of the state, we call it homogeneous, as it models every individual player using a single control $\alpha$. Otherwise, we call it heterogeneous. In the homogeneous case, we do not need to keep track of the flow of the distribution of controls. Moreover, in the heterogeneous case, one can represent the flow of the population $\mu^{\Xi}$ using a single relaxed control instead of a distribution of controls. See [19] for the details.
Introduce the flow of the distribution for the representative player;
$\mathbb{P}^{t,\Xi;x,\alpha}(X_{s+1}=dy|X_{s}=\tilde{x})=p(s,\tilde{x},\mu_{s}^{\Xi},\alpha(s,\tilde{x},\mu_{s}^{\Xi});dy)\ \ \forall t\leq s,\ \ \mathbb{P}^{t,\Xi;x,\alpha}(X_{t}=x)=1$ (3.2)
where $X$ is the canonical process. Notice that the player is observing the distribution of the population $\mu^{\Xi}$, given the initial data $\Xi\in{\cal P}(\mathbb{S}_{t}\times{\cal A})$.
[…]
$J(t,\Xi;x,\alpha):=\mathbb{E}^{t,\Xi;x,\alpha}\Big{[}\phi(X_{t+T},\mu_{t+T}^{\Xi})+\sum_{s=t}^{t+T-1}F(s,X_{s},\mu_{s}^{\Xi},\alpha(s,X_{s},\mu_{s}^{\Xi}))\Big{]},\quad\mathbb{E}^{\cdot}:=\mathbb{E}^{\mathbb{P}^{\cdot}}$
Set $\mathscr{M}_{J}$ as the space of all such functions $(\mathbb{T}\times{\cal P}(\Omega)\times\Omega\times{\cal A}\to\mathbb{R})$. Let us note that, we are mainly interested in the static $\{0,\dots,T\}$ problem for simplicity. One can dynamically set $\hat{T}=T-t$ (and repeats after $T$) by the learning algorithm to create a dynamic version. Or one might evolve the game indefinitely, keeping the $\hat{T}$ fixed.
$\hat{\pi}:\mathbb{T}\times{\cal P}(\Omega)\times\Omega\to{\cal P}({\cal A})$
There is no randomness in the value, and hence if the representative player is able to solve for the optimal control, $\hat{\pi}$ becomes deterministic, or in general, takes values on the set of optimal controls. Moreover, induced distribution is simply ${\Upsilon}^{t,\mu;x}(d\alpha):=\hat{\pi}(t,\mu,x)$ and
$\gamma^{t,\mu;x}(da):={\Upsilon}^{t,\mu;x}\Big{(}\alpha\ :\ \alpha(t,x,\mu)=da\Big{)}\in{\cal P}(\mathbb{A})$
Finally, we are ready to introduce the crucial part: observations and the learning algorithm. Let us briefly recap the mean-field framework. We assume that the representative player starts with an initial guess of the population distribution over states and controls, determines the corresponding optimal control, and, relying on the assumption that everyone else is exactly the same, generates further observations using the chosen learning algorithm. To formalize this, let ${\cal E}={\cal P}({\cal P}(\mathbb{S}\times{\cal A}))$ be the space of observables, and let $\mathscr{E}$ denote the space of finite sequences of ${\cal E}$. Recall that the learning algorithm is a mapping $\mathfrak{L}_{\varphi}:\mathscr{E}\to\mathscr{M}_{\varphi}$, where the player’s estimation at age $n$ is ${}^{n}\mathfrak{L}_{\varphi}:=\mathfrak{L}_{\varphi}({}^{n}{\cal O})$, and the player’s prior is ${}^{0}\mathfrak{L}_{\varphi}$. Here, ${\cal O}:\Omega^{u}\times\mathbb{N}\to\mathscr{E}$ represents the increasing sequence of observations. The set $\Omega^{u}$ corresponds to a potential random choice made by the player during the search for optimal control. To provide intuition, we will construct a simple but explicit learning algorithm for $\mathfrak{L}_{\Gamma}$:
Suppose that the current distribution of the population at time 00 is $\mu\in{\cal P}(\mathbb{S}_{0})$ and is fixed as given. We, as the representative player, start with an initial guess ${}^{0}{\cal O}={}^{0}\Xi=\delta_{(\mu,\delta_{{}^{0}\alpha})}$ for some ${}^{0}\alpha$. That is, our initial observation is $\delta_{{}^{0}\Xi}$. Then, we determine the population flow $\mu^{{}^{0}\Xi}$ using (3.1), our flow $\mathbb{P}^{t,{}^{0}\Xi;x,\tilde{\alpha}}$ using (3.2), and solve the optimization problem to find an optimal control ${}^{1}\alpha\in\hbox{\rm supp$\,$}{\Upsilon}^{0,\mu;x}$.
Now, following the fixed point idea, we learned that if the population is using ${}^{0}\alpha$, it is optimal to use ${}^{1}\alpha$. Since every player is equivalent, we may deduce that the population will use ${}^{1}\alpha$ with some probability $c$, and use ${}^{0}\alpha$ otherwise. That is, the learning algorithm yields
@@
Notice that, for simplicity, we are assuming a homogeneous population. That is, everyone is assumed to be using a single control. Now, we can repeat the same procedure to find another optimal control under the guess $\hat{\Gamma}=\mathfrak{L}_{\Gamma}(\delta_{(\mu,\delta_{{}^{0}\alpha})})$, denoted as ${}^{2}\alpha$, and so on. In general, our naive learning algorithm depends only on the last observation, defined as
[…]
where we took ${}^{n}{\cal O}\in{\cal P}({\cal P}(\mathbb{S}_{0}\times{\cal A}))$ as a single observation rather than a sequence to simplify notation, and ${}^{n+1}\alpha$ is an optimal control under ${}^{n}{\cal O}$. We are assuming that $\hat{\pi}$ yields exact optimal controls, and we do not explicitly account for the possibility of multiple optimal controls. There might be a deterministic selection, or it could be randomly selected. In the latter case, $\mathbb{P}^{u}$ characterizes this random choice.
@@
[…]
@@
Lastly, we rephrase the definition of uncertain equilibrium and will explicitly compute the equilibrium under this basic learning algorithm in the next section.
@@
We say $(\hat{\Gamma},\hat{\pi})\in(\mathscr{M}_{\Gamma},\mathscr{M}_{\pi})$ is an $(\varepsilon,r,\delta)$-uncertain equilibrium at $(t,x,\mu)\in\mathbb{T}\times\mathbb{S}_{t}\times{\cal P}(\mathbb{S}_{t})$ under the learning algorithm $(\mathfrak{L}_{\Gamma},\mathfrak{L}_{\pi})$ if,
@@
    $(\hat{\Gamma},\hat{\pi})$ is the prior of the player,
@@
    $\sup_{\alpha\in{\cal A}}{}^{n}J(t,\mu;x,\alpha)-\int_{{\cal A}}{}^{n}J(t,\mu;x,\alpha){}^{n}\hat{\pi}(t,\mu,x)(d\alpha)\leq\varepsilon,\quad\forall n\in\mathbb{N}$
  • (iii)
    $\mathbb{P}^{u}\big{(}\operatorname*{liminf}_{n\to\infty}d^{t,\mu;x}({}^{0}{\Upsilon}_{\mathfrak{L}}^{t,\mu;x},{}^{n}{\Upsilon}_{\mathfrak{L}}^{t,\mu;x})>r\big{)}\leq\delta$
    where $d^{t,\mu;x}$ is the metric the player equips the space ${\cal P}({\cal A})$ under the equivalence $=^{t,\mu;x}$.
@@ 3.1 One step stated mean-field game examples
$J(\Xi;\alpha):=\mathbb{E}^{\mathbb{P}^{\Xi,\alpha}}\Big{[}\phi(X_{1},\mu_{1}^{\Xi})+F(\alpha)\Big{]}$
$\text{where}\quad\phi(x,\mu):=2|\mu(1)|^{2}-4{\bf 1}_{\left\{{x=1}\right\}}\mu(1),\ \ \text{and}\ \ F(a)=(|a|^{2}+a)$
@@
Proof  First, let us argue that there exists no standard Nash equilibrium. Main idea is, if the population distribution is symmetric $\mu_{1}^{\Xi}(1)=\mu_{1}^{\Xi}(0)=1/2$, then the optimal actions are $\{0,1\}$. Whenever $\mu_{1}^{\Xi}(1)>1/2$, optimal action becomes 00 and otherwise $1$. That is, every player tries to stay away from the majority and there cannot be a deterministic fixed point.
@@
$J(a,\tilde{a}):=J(\Xi;\alpha)$ $=2|\mu^{a}(1)|^{2}-4\mu^{a}(1)\mathbb{P}^{\Xi,\alpha}(X_{1}=1)+|\tilde{a}|^{2}+\tilde{a}$
@@
and in this case, again from [19], equilibrium means that every action in the support of $\Xi_{0}$ is optimal. It is easy to check that if $\mu^{\Xi}(1)\neq 1/2$, then the optimal action is either 00 or $1$ and there cannot be any equilibrium. If $\mu^{\Xi}(1)=1/2$, then both 00 and $1$ are optimal. Thus, $\Xi_{0}=\tfrac{1}{2}(\delta_{0}+\delta_{1})$ corresponds to the relaxed equilibrium, since any action in the support is optimal.
@@
which is exactly as before a quadratic polynomial in $\tilde{a}$, hence optimal value occurs at either 00 or $1$. Therefore, the algorithm (3.4) will quickly converge to a distribution having $\delta_{0},\delta_{1}\in{\cal P}(\mathbb{A})$ in its support, and the contribution from the initial guess will diminish with the factor $(1-c)^{n}$. Moreover, as the optimal $\tilde{a}$ becomes 00 or $1$ depending on the estimated average of the population similarly as before, the algorithm will oscillate around $\frac{1}{2}(\delta_{\delta_{0}}+\delta_{\delta_{1}})$. Note that, by adjusting the constant parameter $c$ in (3.4), one can achieve exact convergence too [^10] Let us briefly take attention to the importance of the design of the learning algorithm, even for this simple setting. If the parameter $c$ diminishes very fast with $n$, then one can conclude either $\delta_{0}$ or $\delta_{1}$ is optimal depending on the initial condition.. Here, since $\hat{\pi}$ is computed exactly as either $\delta_{0}$ or $\delta_{1}$, we also observe that the induced distribution oscillates in between them infinitely often.
@@
[…]
Proof  The essence of this example is similar to that in Example 1. It is immediately clear that there exists no relaxed equilibrium, as the optimal action is either $\delta_{0}$ or $\delta_{1}$ under any value of $\bar{\mu}_{1}^{\Xi}$, and neither constitutes an equilibrium.
@@ 4 Reinforcement Learning
We now specialize the framework of Section 2 to a single player, reducing the problem to a control setting. The main objective is to show that familiar reinforcement learning (RL) methods naturally fall within the scope of this framework. Many RL designs can be viewed as instantiations of the player defined in Definition 1.  [^11] In evolutionary approaches, it is more natural to work with a population of players, as in Section 2. However, the evolutionary mechanisms needed are not formalized in this work.
The simplest illustration is vanilla $Q$-learning. The agent maintains a single space of estimations

### v4

@@ 4 Stated Mean-Field Games
[…]
@@
$\Xi_{s+1}(dy,d\alpha)=\int_{\mathbb{S}_{t}}p(s,x,\mu_{s}^{\Xi},\alpha(s,x,\mu_{s}^{\Xi});dy)d\Xi_{s}(x,d\alpha),\ \forall t\leq s,$ (4.1)
where $\mu_{s}^{\Xi}:=\Xi_{s}(\cdot,{\cal A})$. Note that $\mu^{\Xi}$ corresponds to (3.3) for the population. If the second marginal of $\Xi$ is a Dirac measure $\delta_{\alpha}$ independent of the state, we call it homogeneous, as it models every individual player using a single control $\alpha$. Otherwise, we call it heterogeneous. In the homogeneous case, we do not need to keep track of the flow of the distribution of controls. Moreover, in the heterogeneous case, one can represent the flow of the population $\mu^{\Xi}$ using a single relaxed control instead of a distribution of controls. See [40] for the details.
Next, introduce the flow of the distribution for the representative player;
$\mathbb{P}^{t,\Xi;x,\alpha}(X_{s+1}=dy|X_{s}=\tilde{x})=p(s,\tilde{x},\mu_{s}^{\Xi},\alpha(s,\tilde{x},\mu_{s}^{\Xi});dy)\ \ \forall t\leq s,\ \$(4.2)
with initial condition $\mathbb{P}^{t,\Xi;x,\alpha}(X_{t}=x)=1$ where $X$ is the canonical process. Notice that the player is observing the distribution of the population $\mu^{\Xi}$, given the initial data $\Xi\in{\cal P}(\mathbb{S}_{t}\times{\cal A})$.
$J(t,\mu;x,\alpha)=\int_{{\cal P}(\mathbb{S}_{t}\times{\cal A})}J(t,\Xi;x,\alpha)d\hat{\Gamma}_{t}(\Xi),\ \ \text{where}\ \ \hat{\Gamma}_{t}((\mu,{\cal P}({\cal A})))=1,\text{ and}$ (4.3)
$J(t,\Xi;x,\alpha)=\mathbb{E}^{t,\Xi;x,\alpha}\Big[\phi(X_{t+T},\mu_{t+T}^{\Xi})+\sum_{s=t}^{t+T-1}F(s,X_{s},\mu_{s}^{\Xi},\alpha(s,X_{s},\mu_{s}^{\Xi}))\Big],$
Set $\mathscr{M}_{J}$ as the space of all such functions $(\mathbb{T}\times{\cal P}(\Omega)\times\Omega\times{\cal A}\to\mathbb{R})$. Let us note that, we are mainly interested in the static $\{0,\dots,T\}$ problem for simplicity. One can dynamically set $\hat{T}=T-t$ (and repeats after $T$) by the learning algorithm to create a dynamic version. Or one might evolve the game indefinitely, keeping the $\hat{T}$ fixed if the structure allows it.
$\hat{\pi}:\mathbb{T}\times{\cal P}(\Omega)\times\Omega\to{\cal A}$
There is no randomness in the value, and assuming that the value is sufficiently representative, we don’t further impose randomness in the policy. Thus, given that representative player is able to solve for the optimal control, $\hat{\pi}$ becomes deterministic, taking values on the set of optimal controls. Moreover, behavior is simply ${\Upsilon}(\hat{\pi})=((t,\mu;x)\mapsto\hat{\pi}(t,\mu;x))$. Due to the time-consistency, optimal policy determined at an initial condition stays optimal, but we will rely only on $\hat{\pi}$ to generate observations rather than the behavior.
Let us briefly recap the mean-field framework. We assume that the representative player starts with an initial guess of the population distribution over states and controls, determines the corresponding optimal control, and, relying on the assumption that everyone else is exactly the same, generates further observations using the chosen learning algorithm. Equivalently, one can say that there are infinitely many such players actually playing the game and observing the distribution of players, however, our framing is more consistent with applications. For example, let us consider a portfolio liquidation problem that someone faces in a financial market. Instead of solving an optimization problem without acknowledging that there are other players also facing a similar problem, as a first order approach without actually having information about other players, the player can model there is a distribution of others facing exactly the same problem.
@@
Suppose that the current distribution of the population at time 0 is $\mu\in{\cal P}(\mathbb{S}_{0})$ and is fixed as given. We, as the representative player, start with an initial guess $^{0}{\cal O}={}^{0}\Xi=\delta_{(\mu,\delta_{{}^{0}\alpha})}$ for some $^{0}\alpha$. That is, our initial observation is $\delta_{{}^{0}\Xi}$. Then, we determine the population flow $\mu^{{}^{0}\Xi}$ using (4.1), our flow $\mathbb{P}^{t,{}^{0}\Xi;x,\tilde{\alpha}}$ using (4.2), and solve the optimization problem to find an optimal control $^{1}\alpha$.
Now, following the fixed point idea, we learned that if the population is using $^{0}\alpha$, it is optimal to use $^{1}\alpha$. Since every player is equivalent, we may deduce that the population will use $^{1}\alpha$ with some probability $c$, and use $^{0}\alpha$ otherwise. That is, we set the learning algorithm as
@@
Notice that, for simplicity, we are assuming a homogeneous population. That is, everyone is assumed to be using a single control. Once can, for example, formulate it as a portion of the population will use $^{1}\alpha$. Next, we can repeat the same procedure to find another optimal control under the guess $\hat{\Gamma}=\mathfrak{L}_{\Gamma}(\delta_{(\mu,\delta_{{}^{0}\alpha})})$, denoted as $^{2}\alpha$, and so on. In general, our naive learning algorithm depends only on the last observation, defined as
$\mathfrak{L}_{\Gamma}\big({}^{n}{\cal O}\big):=c\ \delta_{(\mu,\delta_{{}^{n+1}\alpha})}+(1-c)\ \ {}^{n}{\cal O},\ \ 0\leq c\leq 1$ (4.4)
where we took $^{n}{\cal O}\in{\cal P}({\cal P}(\mathbb{S}_{0}\times{\cal A}))$ as a single observation rather than a sequence to simplify notation, and $^{n+1}\alpha$ is an optimal control under $^{n}{\cal O}$.
@@
$J^{\text{fictitious}}(t,\mu;x,\alpha):=J\bigg(t,\int_{{\cal P}(\mathbb{S}_{t}\times{\cal A})}\Xi d\hat{\Gamma}(\Xi);x,\alpha\bigg)$
@@
Lastly, we rephrase the definition of uncertain equilibrium with notations as in Section 2, and we explicitly compute the equilibrium under this basic learning algorithm in the next section.
@@
We say that a player $({\cal O},\mathfrak{L}_{\Gamma},\mathfrak{L}_{\pi},{\Upsilon})$ admit $^{*}{\Upsilon}\in\mathscr{M}_{\Upsilon}$ as an $(\varepsilon,\delta)$-uncertain equilibrium under the metric $d$ on $\mathscr{M}_{\Upsilon}$ at $(t,x,\mu)\in\mathbb{T}\times\mathbb{S}_{t}\times{\cal P}(\mathbb{S}_{t})$, if for any prior $^{0}\alpha\in{\cal A}$,
@@
    $\operatorname*{limsup}_{n\to\infty}\big(\sup_{\tilde{\alpha}\in{\cal A}}{}^{n}J(t,\mu;x,\tilde{\alpha})-{}^{n}J(\cdot,{}^{n}\hat{\pi})(t,\mu;x)\big)\leq\varepsilon$
@@
    $\operatorname*{liminf}_{n\to\infty}d({}^{*}{\Upsilon},{}^{n}{\Upsilon})\leq\delta$
@@ 4.1 One step stated mean-field game examples
$J(\Xi;\alpha)=\mathbb{E}^{\mathbb{P}^{\Xi,\alpha}}\Big[\phi(X_{1},\mu_{1}^{\Xi})+F(\alpha)\Big]$
$\text{where}\quad\phi(x,\mu)=2|\mu(1)|^{2}-4{\bf 1}_{\left\{{x=1}\right\}}\mu(1),\ \ \text{and}\ \ F(a)=(|a|^{2}+a)$
@@
Proof   First, let us argue that there exists no standard Nash equilibrium. Main idea is, if the population distribution is symmetric $\mu_{1}^{\Xi}(1)=\mu_{1}^{\Xi}(0)=1/2$, then the optimal actions are $\{0,1\}$. Whenever $\mu_{1}^{\Xi}(1)>1/2$, optimal action becomes 0 and otherwise $1$. That is, every player tries to stay away from the majority and there cannot be a deterministic fixed point.
@@
$J(a,\tilde{a})=J(\Xi;\alpha)$ $=2|\mu^{a}(1)|^{2}-4\mu^{a}(1)\mathbb{P}^{\Xi,\alpha}(X_{1}=1)+|\tilde{a}|^{2}+\tilde{a}$
@@
and in this case, again from [40], equilibrium means that every action in the support of $\Xi_{0}$ is optimal. It is easy to check that if $\mu^{\Xi}(1)\neq 1/2$, then the optimal action is either 0 or $1$ and there cannot be any equilibrium. If $\mu^{\Xi}(1)=1/2$, then both 0 and $1$ are optimal. Thus, $\Xi_{0}=\tfrac{1}{2}(\delta_{0}+\delta_{1})$ corresponds to the relaxed equilibrium, since any action in the support is optimal.
@@
$=2\sum_{i}c_{i}|a_{i}|^{2}-4\tilde{a}\sum_{i}c_{i}a_{i}+|\tilde{a}|^{2}+\tilde{a}$
which is exactly as before a quadratic polynomial in $\tilde{a}$, hence optimal value occurs at either 0 or $1$. Therefore, the algorithm (4.4) will quickly converge to a distribution having $\delta_{0},\delta_{1}\in{\cal P}(\mathbb{A})$ in its support, and the contribution from the initial guess will diminish with the factor $(1-c)^{n}$. Moreover, as the optimal $\tilde{a}$ becomes 0 or $1$ depending on the estimated average of the population similarly as before, the algorithm will oscillate around $\frac{1}{2}(\delta_{\delta_{0}}+\delta_{\delta_{1}})$. Note that, by adjusting the constant parameter $c$ in (4.4), one can achieve exact convergence too [^9] Let us briefly take attention to the importance of the design of the learning algorithm, even for this simple setting. If the parameter $c$ diminishes very fast with $n$, then one can conclude either $\delta_{0}$ or $\delta_{1}$ is optimal depending on the initial condition.. Here, since $\hat{\pi}$ is computed exactly as either $\delta_{0}$ or $\delta_{1}$, we also observe that the induced distribution oscillates in between them infinitely often.
@@
where $\bar{\mu}^{\Xi}:=\int_{[0,1]}x\,d\mu^{\Xi}$. Whereas no relaxed equilibrium exists, the learning algorithm described in (4.4) again oscillates around $\frac{1}{2}(\delta_{\delta_{0}}+\delta_{\delta_{1}})\in{\cal P}({\cal P}(\mathbb{A}))$, and induces an action distribution $\delta_{0}$ or $\delta_{1}$ infinitely often.
Proof   The essence of this example is similar to that in Example 1. It is clear that there exists no relaxed equilibrium, as the optimal action is either $\delta_{0}$ or $\delta_{1}$ under any value of $\bar{\mu}_{1}^{\Xi}$, and neither constitutes an equilibrium.
@@ 5 Reinforcement Learning
In this section, we review several core reinforcement-learning methods using the terminology developed within this framework. Our goal is not to survey the extensive literature, but rather to emphasize the complexity inherent in player design. In the following subsection, we consider the framework of Section 3 in the single-player setting to illustrate a learning method that does not primarily rely on Bellman-type updates. In the following subsection, we review multi-agent reinforcement learning methods.
Let us first review some main categories before detailing them further.
  • (i)
    Value-based design: The player $({\cal O},\mathfrak{L}_{\varphi},{\Upsilon})$ is defined so that the learning algorithms take the form $\mathfrak{L}_{\varphi}:\mathscr{E}\times\mathscr{M}_{\varphi}\times\mathscr{M}_{\Upsilon}\to\mathscr{M}_{\varphi}$, and the behavior is given by ${\Upsilon}:\mathscr{M}_{\varphi}\to\mathscr{M}_{\Upsilon}$. A simple example is $\mathscr{M}_{\varphi}:=\{\mathbb{S}\times\mathbb{A}\to\mathbb{R}\}$ and $\mathscr{M}_{\Upsilon}:=\{\mathbb{S}\to\mathbb{A}\}$, where $\Upsilon$} yields the maximizing argument over $\mathbb{A}$.
  • (ii)
    Policy-based design: The player $({\cal O},{\Upsilon})$ does not rely on an estimate but instead trains the behavior directly. Here, the behavior takes the form ${\Upsilon}:\mathscr{E}\times\mathscr{M}_{\Upsilon}\to\mathscr{M}_{\Upsilon}$.
  • (iii)
    Actor-critic design: The player $({\cal O},\mathfrak{L}_{\varphi},{\Upsilon})$ is defined so that the behavior takes the form ${\Upsilon}:\mathscr{E}\times\mathscr{M}_{\varphi}\times\mathscr{M}_{\Upsilon}\to\mathscr{M}_{\Upsilon}$. Here, estimations are trained similarly to value-based designs and are used to enhance the training of the policy-based behavior. The learning algorithms might also, and typically do, take the form $\mathfrak{L}_{\varphi}:\mathscr{E}\times\mathscr{M}_{\varphi}\times\mathscr{M}_{\Upsilon}\to\mathscr{M}_{\varphi}$ if they aim to learn the value estimate corresponding to the current behavior.
The simplest illustration is vanilla $Q$-learning. The agent maintains a single estimation space